Identify a type of company in your pathway that might purchase fixed assets A company in my pathway that might purchase fixed assets is Bank of America.

List 5 fixed assets that they might purchase to run their business.

1. Furniture (filing cabinets, sofas, desks, chairs etc.)
2. ATM Machines
3. Computer Equipment (routers, servers.)
4. Computer Software (only the most expensive types)
5. Office Equipment (photocopiers, fax machines, postage meter etc.)

I answered the top two questions, but have to answer the others listed below.


Select one depreciable fixed asset. Based on research suggest what the cost, residual value and estimated life might be for that fixed asset. (please list the name of the asset you would be calculating for)

Using your assumptions above, calculate:

Straight-line depreciation and book value for each of the first two years

Declining Balance depreciation and book value for each of the first two years

Units of Production depreciation (make assumptions about the first two year’s use), and book value for each of the first two years.
Answer #1

Select one depreciable fixed asset.

Name- Office Equipment -Photocopier

Cost = $ 800

Residual value = $ 50

Useful life = 5 years

Units of Production = 5,000 copies per month i.e 300,000 copies in 5 years of useful life ( 5,000 * 60 months)

a) Annual depreciation as per Straight-line method = (Cost - Residual value) / useful life in years

= ( 800 - 50) / 5 = $ 150

Depreciation for 1st and 2nd year will be $ 150 each year

Book value at the end of 1st year = $ 800 - 150 = $ 650

Book value at the end of 2nd year = $ 650 - 150 = $ 500

b) Declining Balance rate = 1 / useful life in years * 100 = 1 / 5 * 100= 20%

Depreciation for 1st year = $ 800 * 20% = $ 160

Book value at the end of 1st year = $ 800 - 160 = $ 640

Depreciation for 2nd year = $ 640 * 20% = $ 128

Book value at the end of 2nd year = $ 640 - 128 = $ 512

c) Per unit Depreciation as per Units of Production method = (Cost - Residual value) / Total units to be produced during the useful life

= (800 - 50) / 300,000

= 0.0025 per copy

Depreciation for 1st year = (5,000 * 12) * 0.0025 = $ 150 (5000 copies in each month for 1st year)

Depreciation for 2nd year = (5,500 * 12) * 0.0025 = $ 165 (5500 copies in each month for 2nd year)

Book value at the end of 1st year = $ 800 - 150 = $ 650

Book value at the end of 2nd year = $ 650 - 165 = $ 485